Operations Report- December 20th

Hello! We had a great first day. Thank you for your support!

SOL: 1

Name of person filing report: Pavithra Ravi

Non-nominal systems: Toilet at HAB

Notes on non-nominal systems: Toilet has a leak. A plastic container
is being used to collect the water, which is then sent down the
toilet.

Generator:

Hours run: Information provided by Outpost

From what time last night: Information provided by Outpost

To what time this morning: 7:20 AM

List any additional daytime hours when the generator was run: None

Solar— SOC % (Before generator is run at night): Information provided by Outpost

Station Propane Reading – % Information provided by Outpost

Ethanol Free Gasoline: Not in use

Water (loft tank): 55 gallons

Water Meter: 1539024 units

Water (static tank): approx. 415 gallons

Static to Loft Pump used – yes. Refilled at 7:21 PM from 21 to 55 Gallons.

Water in GreenHab: approx. 240.5 gallons

Water in ScienceDome: N/A

Toilet tank emptied: yes, emptied at 11:45 AM

Deimos rover used: No

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Spirit rover used: Yes

Hours (before EVA 1): 146.3

Beginning charge (Before EVA 1): 100%

Ending charge (On return from EVA 1, before recharging): 99%

Rover was charged before EVA 2.

Hours (before EVA 2): 146.4

Beginning charge (Before EVA 2): 100%

Ending charge (On return from EVA 2, before recharging): 100%

Currently charging: Yes

Opportunity rover used: Yes, for EVA 2

Hours: 76.3

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Yes, for EVA 1

Hours: 154.6

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: None

ATV’s Used (Honda, 300, 350.1, 350.2, 350.3): None

Reason for use: N/A

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were used today: N/A

Notes on ATVs: N/A

HabCar used and why, where? No

CrewCar used and why, where? No

General notes and comments: None

Summary of internet: MOS-MOCC wifi network is no longer available.
Only the regular Mars Hab wifi is in use.

Summary of suits and radios: Nominal

Summary of Hab operations: Nominal. Small concern: the window above
the couch in the Hab does not fit and is letting air in.

Summary of GreenHab operations: Nominal

Summary of ScienceDome operations: Nominal

Summary of RAM operations: Nominal

Summary of any observatory issues: Problems with the dome movement
have been reported to the astronomy support team.

Summary of health and safety issues: Nominal

Questions, concerns and requests to Mission Support: None